OpenStudy (anonymous):

whats an equation that has variables on both sides and the solution is -3

OpenStudy (texaschic101):

3x - 1 = 2x -4 -- subtract 2x from both sides 3x - 2x - 1 = -4 -- add 1 to both sides 3x - 2x = -4 + 1 -- combine like terms x = -3